# Strange Stars Admixed with Mirror Dark Matter: Confronting Observations of XTE J1814-338

### v1

@@ 3.2 Parameter Space and Constraints
We have further explored the parameter space of the SQM model by applying four observational constraints (e.g., Schaab et al., 1997; Weissenborn et al., 2011; Zhou et al., 2018; Yang et al., 2020, 2021a, 2021b, 2023; Yang & Pi, 2024). These constraints are essential to ensure that the model can explain the observed properties of NSs and SSs:
@@
The energy per baryon of SQM must be lower than that of the most stable atomic nucleus, 56Fe, which has an energy per baryon of $E/A\sim 930$ MeV (Witten, 1984; Zhou et al., 2018). This condition ensures that SQM is absolutely stable. The parameter region that satisfies this constraint is located below the 3-flavor line in Fig. 2.
@@
Non-strange quark matter (i.e., two-flavor quark matter consisting of only $u$ and $d$ quarks) must have an energy per baryon higher than 56Fe, plus a 4 MeV surface correction (Farhi & Jaffe, 1984; Zhou et al., 2018; Yang et al., 2020). This constraint ensures that ordinary nuclear matter remains stable and does not dissolve into quark matter. The region satisfying this constraint lies above the 2-flavor line in Fig. 2.
@@
Figure 3: The mass-radius relation ($R_{Q}$ represents the radius of the SQM) of MDM-admixed SSs is shown for $\alpha_{S}=0.6$ and $B^{1/4}=135$ MeV. The magenta line corresponds to $f_{D}=50\%$. From right to left, the black lines represent $f_{D}=0\%$, 10%, 20%, 30%, and 40%, while the grey lines represent $f_{D}=60\%$, 70%, 80%, and 90%. The solid and dashed blue lines indicate $f_{D}=71.8\%$ and $f_{D}=82.2\%$, respectively. The data points and shaded regions are the same as those in figure 1.
@@
The third constraint is derived from the mass and radius measurements of PSR J0437-4715 (Choudhury et al., 2024). The region that satisfies this constraint is located above the blue line in Fig. 2 (note that the blue line corresponds to $M=1.38\,M_{\odot}$ and $R=12.31$ km, which is the coordinate of the lower-right point of the green rectangle in figure 1).
@@
The fourth constraint is based on the mass and radius estimates of PSR J0740+6620 (Salmi et al., 2024). The parameter region satisfying this constraint is below the red line in Fig. 2 (the red line corresponds to $M=2.00\,M_{\odot}$ and $R=11.61$ km, which is the coordinate of the lower-left point of the cyan rectangle in figure 1).
[…]
@@
For comparison, the lines for three different values of $M_{\rm max}$ ($M_{\rm max}$ is the maximum mass of SSs without MDM) and $\Lambda(1.4)=580$ [$\Lambda(1.4)$ is the dimensionless tidal deformability of a 1.4 $M_{\odot}$ pure SS] are also shown in Fig. 2. The regions below the grey solid, dashed and dotted lines correspond to the parameters that satisfy $M_{\rm max}>2.0\,M_{\odot}$, $M_{\rm max}>2.2\,M_{\odot}$, and $M_{\rm max}>2.4\,M_{\odot}$, respectively. The region above the green line fulfills the $\Lambda(1.4)<580$ constraint from the observation of GW170817 (Abbott et al., 2018).
3.3 Inclusion of MDM
@@
This two-fluid behavior is crucial for explaining the mass and radius observations of XTE J1814-338 (Kini et al., 2024). Fig. 3 shows that the data for XTE J1814-338 can be explained if the star is an MSS with an ordinary SQM core, where the mass fraction of MDM falls between $f_{D}=71.8\%$ and $f_{D}=82.2\%$. This is consistent with previous results, suggesting that XTE J1814-338 belongs to a class of stars that include both SQM and MDM.
4 Testing SSs with MDM via Multi-Messenger Astronomy
@@
The predictions of our SS and MDM admixed model can be rigorously tested through various observational approaches. Gravitational wave detections provide a powerful method for probing the internal structure of compact stars. Events like GW170817, which have constrained tidal deformability, set the stage for testing our model. We anticipate that the SS+MDM configuration will produce distinctive tidal deformability signatures in future observations of binary star mergers. These signatures could serve as a decisive test of the model, particularly when analyzed with current detectors like LIGO and Virgo, and enhanced by next-generation observatories.
X-ray timing observations, particularly from missions like NICER, are another key avenue for testing the model. Precise measurements of NS masses and radii, such as those for PSR J0740+6620 and XTE J1814-338, offer direct comparisons with the compactness predictions of the SS+MDM model. By comparing the observed data with the predictions of our model, we can assess the validity of the model and make necessary refinements.
@@
Thermal emission studies also offer valuable insights for observational testing. The presence of an MDM halo is expected to influence the thermal evolution and cooling behavior of compact stars. By analyzing deviations from standard cooling curves, we can identify potential signatures pointing to the involvement of MDM.
Ultimately, multi-messenger astronomy—which combines gravitational wave and electromagnetic observations—will be crucial in thoroughly testing the SS+MDM hypothesis. By integrating data from multiple observational channels, we can gain a comprehensive understanding of the model and distinguish it from other compact star models, such as those involving only SQM or purely nuclear matter.
5 Summary
@@
First, we successfully explained the mass and radius observations of PSR J0740+6620, PSR J0030+0451, PSR J0437-4715, and the central compact object within the supernova remnant HESS J1731-347. These objects were modeled as pure SSs without the need for additional dark matter components. The results show that SSs alone can account for the majority of observational data, further supporting the hypothesis that SQM may exist as the true ground state of baryonic matter. The results are consistent with previous studies and confirm that SSs provide a plausible explanation for these NS observations.
Horvath et al. (2023) examined XMMU J173203.3-344518 in the remnant HESS J1731-347, proposing it as an SS due to its low mass and small radius, which challenge traditional NS models. Their analysis, based on quark matter equations of state and color-flavor-locked phases, supports the idea that SSs could explain both low-mass and heavier compact objects. This aligns with our findings for XTE J1814-338, strengthening the case for SQM in exotic compact stars. Our model, incorporating MDM, complements their work by offering a mechanism to explain unusual compact star properties, expanding the scope of SS candidates.
Recently, Giangrandi et al. (2023) investigated the effects of self-interacting bosonic dark matter on NS properties, using a two-fluid model to describe the dark matter and baryonic matter components. Their findings reveal that the presence of dark matter significantly alters the mass, radius, and tidal deformability of NSs, depending on the distribution of dark matter in the core or halo. These results align with our hypothesis of MDM admixed with SSs, further supporting the idea that dark matter may play a crucial role in shaping the properties of compact stellar objects.
@@
XTE J1814-338 can also be explained by alternative models, such as the boson star model proposed by Pitz and Schaffner-Bielich (Pitz & Schaffner-Bielich, 2024), which describes a core of ordinary nuclear matter, or by compact stars admixed with self-interacting fermionic dark matter, as shown by Liu et al. (2024a). While these models provide alternative explanations, the success of our MDM model highlights the viability and potential necessity of dark matter, particularly in the form of MDM, in explaining such exotic observations.
As the field progresses, developing methods to differentiate between various dark matter candidates, such as boson stars, fermionic dark matter, and MDM, will be essential. Upcoming observations from advanced gravitational wave detectors such as aLIGO, aVirgo, Kagra, the Einstein Telescope (ET), and the Cosmic Explorer(CE); and the X-ray missons such as eXTP (Watts et al., 2019), STROBE-X (Ray et al., 2019) and ATHENA (Majczyna et al., 2020), are likely to provide the sensitivity required to detect subtle differences between these models, especially in the context of multi-messenger astronomy.
This work is supported by the National Key R&D Program of China (Grant No. 2021YFA0718504) and the Scientific Research Program of the National Natural Science Foundation of China (NSFC, Grant No. 12033001).

### v4

@@ III.2 Parameter Space and Constraints
We have further explored the parameter space of the SQM model by applying four observational constraints [e.g., 102, 25, 26, 24, 19, 103, 104, 105, 106]. These constraints are essential to ensure that the model can explain the observed properties of NSs and SSs:
@@
The energy per baryon of SQM must be lower than that of the most stable atomic nucleus, 56Fe, which has an energy per baryon of $E/A\sim 930$ MeV [7, 26]. This condition ensures that SQM is absolutely stable. The parameter region that satisfies this constraint is located below the 3-flavor line in Fig. 2.
@@
Non-strange quark matter (i.e., two-flavor quark matter consisting of only $u$ and $d$ quarks) must have an energy per baryon higher than 56Fe, plus a 4 MeV surface correction [8, 26, 24]. This constraint ensures that ordinary nuclear matter remains stable and does not dissolve into quark matter. The region satisfying this constraint lies above the 2-flavor line in Fig. 2.
@@
Figure 3: The mass-radius relation ($R_{Q}$ represents the radius of the SQM) of MDM-admixed SSs for $\alpha_{S}=0.6$ and $B^{1/4}=135$ MeV. The magenta line corresponds to $f_{D}=50\%$. From right to left, the black lines represent $f_{D}=0\%$, 10%, 20%, 30%, and 40%, while the grey lines represent $f_{D}=60\%$, 70%, 80%, and 90%. The solid and dashed blue lines indicate $f_{D}=71.8\%$ and $f_{D}=82.2\%$, respectively. The shaded regions are the same as those in Fig. 1.
Figure 4: The mass-radius relation ($R_{Q}$ represents the radius of the SQM) of MDM-admixed NSs for the KWCZ3 EOS. The values of $f_{D}$ for the magenta line, and the black and grey lines, are the same as those in Fig. 3. The solid and dashed blue lines indicate $f_{D}=72.8\%$ and $f_{D}=85.1\%$, respectively. The shaded regions are identical to those in Fig. 1.
@@
The third constraint is derived from the mass and radius measurements of PSR J0437-4715 [98]. The region that satisfies this constraint is located above the blue line in Fig. 2 (note that the blue line corresponds to $M=1.38\,M_{\odot}$ and $R=12.31$ km, which is the coordinate of the lower-right point of the green rectangle in Fig. 1).
@@
The fourth constraint is based on the mass and radius estimates of PSR J0740+6620 [96]. The parameter region satisfying this constraint is below the red line in Fig. 2 (the red line corresponds to $M=2.00\,M_{\odot}$ and $R=11.61$ km, which is the coordinate of the lower-left point of the cyan rectangle in Fig. 1).
Combining all four constraints, the allowed parameter space is confined to the yellow-shaded area in Fig. 2. Specifically, the parameters of the SQM model are limited to 131.5 MeV $\leq B^{1/4}\leq$ 140.8 MeV (it is slightly different from the case of $\alpha_{S}=0.6$ because different values of $\alpha_{S}$ are involved here) and 0.18 $\leq\alpha_{S}\leq$ 0.86. It is important to note that this parameter space is derived under the assumption that both PSR J0437-4715 and PSR J0740+6620 are pure SSs without MDM. If MDM were included, the results would likely differ, potentially broadening the parameter space.
@@
For comparison, the lines for three different values of $M_{\rm max}$ ($M_{\rm max}$ is the maximum mass of SSs without MDM) and $\Lambda(1.4)=580$ [$\Lambda(1.4)$ is the dimensionless tidal deformability of a 1.4 $M_{\odot}$ pure SS] are also shown in Fig. 2. The regions below the grey solid, dashed and dotted lines correspond to the parameters that satisfy $M_{\rm max}>2.0\,M_{\odot}$, $M_{\rm max}>2.2\,M_{\odot}$, and $M_{\rm max}>2.4\,M_{\odot}$, respectively. The region above the green line fulfills the $\Lambda(1.4)<580$ constraint from the observation of GW170817 [15].
The constraints outlined above provide a comprehensive framework for understanding the properties of SSs under the assumption of pure SQM. However, these models do not account for the potential influence of additional components, such as mirror dark matter (MDM). Incorporating MDM into the equation of state introduces new degrees of freedom, significantly altering the mass-radius relationship and broadening the parameter space for SSs. To investigate these effects, we extend our analysis in the next section to include MDM, exploring its impact on the structure and observational properties of compact stars.
III.3 Inclusion of Mirror Dark Matter (MDM)
@@
This two-fluid behavior is crucial for explaining the mass and radius observations of XTE J1814-338 [11]. Figure 3 shows that the data for XTE J1814-338 can be explained if the star is an MSS with an ordinary SQM core, where the mass fraction of MDM falls between $f_{D}=71.8\%$ and $f_{D}=82.2\%$. This is consistent with previous results, suggesting that XTE J1814-338 belongs to a class of stars that include both SQM and MDM.
Figure 4 illustrates the mass-radius relationship for MDM-admixed neutron stars, modeled using the KWCZ3 EOS (for the case of $C_{\sigma}=14$, $L=40$) [101]. The dependence of the mass-radius curves on the MDM mass fraction ($f_{D}$) is evident, with higher $f_{D}$ values producing more compact configurations. The data for XTE J1814-338, represented by the red shaded region, is explained within a specific $f_{D}$ range of 72.8% to 85.1%. This alignment highlights the potential role of MDM in shaping the structural properties of compact stars. Unlike the MDM-admixed strange stars shown in Fig. 3, the neutron star model in Fig. 4 shows subtle shifts in the mass-radius curves, particularly at higher $f_{D}$. The solid and dashed blue lines denote the $f_{D}$ thresholds where the radius of the neutron star’s baryonic component becomes dominated by the MDM envelope, signifying a transition to a two-layer structure. The inclusion of MDM alters the tidal deformability and compactness of neutron stars, making these models distinguishable from traditional neutron stars. For XTE J1814-338, the high $f_{D}$ values required to explain its observations suggest a substantial dark matter admixture, underscoring the critical role of multi-messenger astrophysical observations.
IV Considerations and Future Perspectives
The predictions of our models of SSs admixed with MDM can be thoroughly tested through multi-messenger astronomy, which combines gravitational wave, X-ray, and thermal emission observations. These observations offer unique opportunities to validate the presence of exotic matter compositions in compact stars, such as XTE J1814-338. In the following, we explore how gravitational wave signatures, precise mass-radius measurements from X-ray timing, and thermal emission studies can be used to test and refine these predictions.
@@ IV Considerations and Future Perspectives
Gravitational wave detections provide a powerful method to probe the internal structure of compact stars. Events such as GW170817 [14] have constrained tidal deformability, setting the stage for testing our model. The SS+MDM configuration is predicted to exhibit distinctive tidal deformability signatures in future observations of binary star mergers. These signatures, observable with advanced detectors like LIGO, Virgo, Kagra, and next-generation observatories (e.g., Einstein Telescope and Cosmic Explorer), could serve as decisive evidence supporting the model.
X-ray timing observations from missions like NICER are crucial for assessing the SS+MDM model. Precise mass and radius measurements, such as those for PSR J0740+6620 and XTE J1814-338, allow direct comparisons with our predictions of compactness and the effects of MDM admixture. Incorporating improved measurements and reducing observational uncertainties will help refine parameter estimates such as the bag constant $B$ and MDM fraction $f_{D}$.
@@
The presence of an MDM halo is expected to influence the thermal evolution and cooling behavior of compact stars. Deviations from standard cooling curves could provide additional observational signatures pointing to MDM. This avenue offers complementary insights and underscores the importance of thermal emission studies alongside mass-radius observations.
While our model provides a compelling explanation for XTE J1814-338, alternative frameworks—such as boson stars [86, 13] or fermionic dark matter models [44, 107]—also predict compact objects with unusual mass-radius properties. Boson stars may involve new scalar particles beyond the Standard Model, which could introduce novel physics but also pose challenges for physical interpretation. Fermionic dark matter stars rely on self-interacting fermions to achieve compactness, a feature that invites further exploration. In contrast, the SS+MDM framework builds upon known physical principles and extends the strange star hypothesis in a natural way. However, like all models, the SS+MDM framework also requires further validation and refinement. Future observations, including measurements of tidal deformability, cooling curves, and gravitational wave signatures, will be critical for testing and distinguishing between these models.
Ongoing and future missions, including eXTP [108, 109], STROBE-X [110] and ATHENA [111], promise enhanced observational precision. These advancements will enable more accurate tests of the SS+MDM framework and help constrain the parameter space of exotic compact objects. Simultaneously, theoretical efforts should explore alternative dark matter candidates and potential direct interactions between SQM and MDM to refine existing models.
Ultimately, multi-messenger astronomy has the potential to provide valuable insights for evaluating the predictions of the SS+MDM model, distinguishing it from competing theories, and deepening our understanding of compact star structures and the role of dark matter in astrophysics.
V Summary
@@
To assess the robustness of our SS+MDM framework, we conducted a sensitivity analysis of key parameters, such as the bag constant B and the MDM mass fraction $f_{D}$. Variations in $B$ within the allowed range (131.5 MeV to 140.8 MeV) were found to significantly affect the predicted radii of SSs, with lower values of $B$ leading to more compact configurations. Similarly, changes in the strong interaction coupling constant $\alpha_{S}$ introduced shifts in the mass-radius curves, emphasizing the importance of constraining these parameters through independent observations. For the MDM component, the mass fraction $f_{D}$ showed the most pronounced influence on the model’s ability to reproduce the observed properties of XTE J1814-338. Stars with $f_{D}$ outside the range of 71.8% to 82.2% were unable to match the measured compactness of XTE J1814-338.
Building on this analysis, we first applied our framework to explain the mass and radius observations of PSR J0740+6620, PSR J0030+0451, PSR J0437-4715, and the central compact object within the supernova remnant HESS J1731-347. These objects were modeled as pure SSs without requiring additional dark matter components. The results demonstrate that standard SSs can account for the majority of the observational data, supporting the hypothesis that SQM may represent the true ground state of baryonic matter. This consistency with previous studies reinforces the plausibility of SSs as an explanation for these NS observations.
Horvath et al. [112] examined XMMU J173203.3-344518 in the remnant HESS J1731-347, proposing it as an SS due to its low mass and small radius, which challenge traditional NS models. Their analysis, based on quark matter equations of state and color-flavor-locked phases, supports the idea that SSs could explain both low-mass and heavier compact objects. This aligns with our findings for XTE J1814-338, strengthening the case for SQM in exotic compact stars. Our model, incorporating MDM, complements their work by offering a mechanism to explain unusual compact star properties, expanding the scope of SS candidates.
Recently, Giangrandi et al. [80] investigated the effects of self-interacting bosonic dark matter on NS properties, using a two-fluid model to describe the dark matter and baryonic matter components. Their findings reveal that the presence of dark matter significantly alters the mass, radius, and tidal deformability of NSs, depending on the distribution of dark matter in the core or halo. These results align with our hypothesis of MDM admixed with SSs, further supporting the idea that dark matter may play a crucial role in shaping the properties of compact stellar objects.
@@
XTE J1814-338 can also be explained by alternative models, such as the boson star model proposed by Pitz and Schaffner-Bielich [13], which describes a core of ordinary nuclear matter, or by compact stars admixed with self-interacting fermionic dark matter, as shown by Liu et al. [44]. While these models provide alternative explanations, the success of our MDM model highlights the viability and potential necessity of dark matter, particularly in the form of MDM, in explaining such exotic observations.
For the explaination of the observation of XTE J1814-338, large values of $f_{D}$ (lager than 70%) should be employed, which could not be obtained from normal accretion processes [74]. However, it is reachable through other mechanisms listed in ref. [74], e.g., it could be realized if a normal SS with a smaller mass mergers with a pure MSS with a larger mass.
As the field progresses, developing methods to differentiate between various dark matter candidates, such as boson stars, fermionic dark matter, and MDM, will be essential. Upcoming observations from advanced gravitational wave detectors such as aLIGO, aVirgo, Kagra, the Einstein Telescope (ET), and the Cosmic Explorer (CE); and the X-ray missons such as eXTP [108, 109], STROBE-X [110] and ATHENA [111], are likely to provide the sensitivity required to detect subtle differences between these models, especially in the context of multi-messenger astronomy.
Acknowledgements.
We thank Sebastian Kubis for kindly providing the KWCZ EOS table. This work is supported by the National Key R&D Program of China (Grant No. 2021YFA0718504) and the Scientific Research Program of the National Natural Science Foundation of China (NSFC, Grant No. 12033001).